# Mitigating the Impact of Labeling Errors on Training via Rockafellian Relaxation

### v1

@@ 3.4.1 Loss-reweighting via Data-Driven Wasserstein Formulation
Given a sample $\{(1,x_{i},y_{i})\}_{i=1}^{N}$, just as in Section 3.2, we can opt not to take as granted the resulting empirical distribution $\mathcal{D}_{N}$ because of the possibility that $|C|$-many have incorrect labels (i.e., $y_{i}\neq\tilde{y}_{i})$. Instead, we will admit alternative distributions obtained by shifting the $\mathcal{D}_{N}$’s probability mass off “corrupted" tuples $(1,x_{i},y_{i})_{i\in C}$ to possibly $(0,x_{i},y_{i})$, $(1,x_{i},\tilde{y}_{i})$, or even some other tuple $(1,x_{j},\tilde{y}_{j})$ with $j\notin C$ for example - equivalently, eliminating, correcting, or replacing the sample, respectively. In order to admit such favorable corrections to $\mathcal{D}_{N},$ we can consider the optimistic [19, 30] data-driven problem
@@
in which for each parameter tuning $\theta$, $v_{N}(\theta)$ measures the expected auxiliary loss with respect to the most favorable distribution within an $\epsilon$ prescribed $W_{1}$ (1- Wasserstein) distance of $\mathcal{D}_{N}$. It turns out that a budgeted deviation of the weights alone (and not the feature-label pairs) can approximate (up to an error diminishing in $N$) $v_{N}(\theta)$. More precisely, we derive the following approximation along similar lines to [30].
@@
$v_{N}^{MIX}(\theta):=\min_{u_{1},\ldots,u_{N}}~{}$ $\sum_{i=1}^{N}(\frac{1}{N}+u_{i})\cdot J(\theta;x_{i},y_{i})+\gamma_{\theta}\sum_{i=1}^{N}|u_{i}|$
@@
In particular, $-\gamma_{\theta}\leq\min_{i}J(\theta;x_{i},y_{i})$, and $\{i:J(\theta;x_{i},y_{i})>\gamma_{\theta}\}$ are all down-weighted to zero, i.e., $u_{i}^{*}=-\frac{1}{N}$ for any $u^{*}$ solving $v_{N}^{MIX}(\theta).$
@@
Hence, in some sense, we can confirm that RRM is an optimistic methodology but that it is less optimistic than the data-driven Wasserstein approach.
3.5 RRM Algorithm
Towards solving problem (3) in the two decisions $\theta$ and $u$, we proceed iteratively with a block-coordinate descent heuristic outlined in Algorithm 1, whereby we update the two separately in cyclical fashion. In other words, we update $\theta$ while holding $u$ fixed, and we update $u$ whilst holding $\theta$ fixed. The update of $\theta$ is an SGD step on a batch of $s-$ many samples. The update of $u$ reduces to a linear program. In light of the discussion in 3.4, we also outline an Adversarial Rockafellian Relaxation method (A-RRM), an execution of RRM that includes a perturbation (parameterized by $\epsilon\geq 0$) to the feature $x$ of a sample $(x,y),$ for the purposes of adversarial training.
@@ 3.5 RRM Algorithm
0:  Perturbation Multiplier $\epsilon\in[0,1]$, Number of epochs $\sigma,$ Batch size $s\geq 1$, learning rate $\eta>0,$ regularization parameter $\gamma>0$, reweighting step $\mu\in(0,1)$.
@@
           $\{(x^{b}_{i},y^{b}_{i})\}_{i=1}^{s}\leftarrow\text{Draw Batch of size $s$ from}\{(x_{i},y_{i})\}_{i=1}^{N}$
@@
     $u\leftarrow\mu u^{*}+(1-\mu)u$
  until Desired Validation Accuracy or Loss
The stepsize parameters $\mu,\eta$ and the regularization parameter $\gamma$ are hyper-parameters that may be tuned, or guided by the general discussions above in Section 3.3.
The RRM algorithm, in which $\epsilon=0,$ is meant for contexts in which only label corruption and no feature corruption occurs. The A-RRM algorithm, for which $\epsilon>0,$ is intended for contexts in which both label and feature corruption is anticipated.
@@ 4 Datasets
We select several datasets to evaluate RRM. In some cases, the selected dataset is nearly pristine. In these cases we perturb the dataset to achieve various types and levels of corruption. Other datasets consist of weakly labeled examples, which we maintain unaltered. The varied data domains and regimes of corruption enable a robust evaluation of RRM.
MNIST [17]: A multi-class classification dataset consisting of 70000 images of digits zero through nine. 60000 digits are set aside for training and 10000 for testing. 0%, 5%, 10%, 20%, and 30% of the training labels are swapped for different, randomly selected digits. The test set labels are unmodified.
Toxic Comments [6]: A multi-label classification problem from JIGSAW that consists of Wikipedia comments labeled by humans for toxic behavior. Comments can be any number (including zero) of six categories: toxic, severe toxic, obscene, threat, insult, and identity hate. We convert this into a binary classification problem by treating the label as either none of the six categories or at least one of the six categories. This dataset is a public dataset used as part of the Kaggle Toxic Comment Classification Challenge.
IMDb [18]: A binary classification dataset consisting of 50000 movie reviews each assigned a positive or negative sentiment label. 25000 reviews are selected randomly for training and the remaining are used for testing. 25%, 30%, 40%, and 45% of the labels of the training reviews are randomly selected and swapped from positive sentiment to negative sentiment, and vice versa, to achieve four training datasets of desired levels of label corruption. The test set labels are unmodified.
Tissue Necrosis: A binary classification dataset consisting of 7874 256x256-pixel hematoxylin and eosin (H&E) stained RGB images derived from [2]. The training dataset consists of 3156 images labeled non-necrotic, as well as 3156 images labeled necrotic. The training images labeled non-necrotic contain no necrosis. However, only 25% of the images labeled necrotic contain necrotic tissue. This type of label error can be expected in cases of weakly-labeled Whole Slide Imagery (WSI). Here, an expert pathologist will provide a slide-level label for a potentially massive slide consisting of gigapixels, but they lack time or resources to provide granular, segmentation-level annotations of the location of the pathology in question. Also, the diseased tissue often occupies a small portion of the WSI, with the remainder consisting of normal tissue. When the gigapixel-sized WSI is subsequently divided into sub-images of manageable size for typical machine-learning workflows, many of the sub-images will contain no disease, but will be assigned the "weak" label chosen by the expert for the WSI. The test dataset consists of 718 necrosis and 781 non-necrosis 256x256-pixel H&E images, which were also derived from [2]. For both the training and test images, [2] provide segmentation-level necrosis annotations, so we are able to ensure a pristine test set, and, in the case of the training set, we were able to identify the corrupted images for the purpose of algorithm evaluation.
@@ 5 Architectures
We do not strive to develop a novel NN architectures capable of defeating current state-of-the-art (SOA) performance in each data domain. Nor do we focus on developing robust architectures as described in [28]. Rather, we select a reasonable NN architecture and measure model performance with and without the application of RRM. This approach enables us to demonstrate the general superiority of RRM under varied data domains and NN architectures. We discuss the underlying NN architectures that we employ in this section.
MNIST: The MNIST dataset has been studied extensively and harnessed to investigate novel machine-learning methods, including CNNs [4]. We adopt a basic CNN architecture with a few convolutional layers. The first layer has a depth of 32, and the next two layers have a depth of 64. Each convolutional layer employs a kernel of size three and the ReLU activation function followed by a max-pooling layer employing a kernal of size 2. The last convolutional layer is connected to a classification head consisting of a 100-unit dense layer with ReLU activation, followed by a 10-unit dense layer with softmax activation. In total, there are 159254 trainable parameters. Categorical cross-entropy is employed for the loss function.
Toxic Comments: We use a simple model with only a single convolutional layer. A pretrained embedding from FastText is first used to map the comments into a 300 dimension embedding space, followed by a single convolutional layer with a kernel size of two with a ReLU activation layer followed by a max-pooling layer. We then apply a 36-unit dense layer, followed by a 6 unit dense layer with sigmoid activation. Binary cross-entropy is used for the loss function.
IMDb: Transformer architectures have achieved SOA performance on the IMDb dataset sentiment analysis task [7, 32]. As such, we a adopt a reasonable transformer architecture to assess RRM. We utilize the DistilBERT [25] architecture with low-rank adaptation (LoRA) [13] for large language models, which reduces the number of trainable weights from 67584004 to 628994. In this manner, we reduce the computational burden, while maintaining excellent sentiment analysis performance. Binary cross-entropy is employed for the loss function.
Tissue Necrosis: Consistent with the computational histopathology literature [21], we employ a convolutional neural network (CNN) architecture for this classification task. In particular, a ResNet-50 architecture with pre-trained ImageNet weights is harnessed. The classification head is removed and replaced with a dense layer of 512 units and ReLU activation function, followed by an output layer with a single unit using a sigmoid activation function. All weights, with the exception of the new classification head are frozen, resulting in 1050114 trainable parameters out of 24637826. Binary cross-entropy is employed for the loss function.
@@ 6 Experiments and Results
In this work, we have discussed errors/perturbations/corruption to features and labels. We now perform experiments to see how RRM performs under one or the other, or both. The MNIST experiments are performed under a setting of both adversarial perturbation, as well as label corruption. The Toxic Comments experiments are performed under settings of label corruption only. All experiments are performed using a combination of GPU resources, both cloud-base, as well as access to an on-premise high-performance computing (HPC) facility. We refer the reader to the Appendix (Sections 6.3 and 6.4) for the experiments on IMDb and Tissue Necrosis.
6.1 MNIST
Twenty percent of the training data is set aside for validation purposes. Using Tensorflow 2.10 [1], 50 iterations of RRM are executed with $\sigma=10$ epochs per iteration for a total of 500 epochs for a given hyperparameter setting. For RRM, the hyperparameter settings of $\mu$ and $\gamma$ at 0.5 and 2.0, respectively, are based on a search to optimize validation set accuracy. For contrast, we perform a comparable 500 epochs using ERM. Both ERM and RRM employ stochastic gradient descent (SGD) with a learning rate ($\eta$) of $0.1$. Each time a batch is drawn, each training image is perturbed using the Fast Gradient Sign Method (FGSM) [11] adversarial attack: $adv_{x}=x+\epsilon\cdot sign(\nabla_{x}J(\theta,x,y))$, where $adv_{x}$ is the resulting perturbed image, $x$ is the original image, $y$ is the image label, $\epsilon$ is a multiplier controlling the magnitude of the image perturbation, $\theta$ are the model parameters, and $J$ is the loss. An $\epsilon=1.0$ is used for all training image perturbations.
For each of the 0%, 5%, 10%, 20%, and 30% training label corruption levels, we compare adversarial training (AT) and adversarial RRM (A-RRM) performance under varios regimes of test set perturbation ($\epsilon_{test}\in{0.0,0.1,0.25,0.5,1.0}$). In Table 1 we show the test set accuracy achieved when validation set accuracy peaks. We can see that training with an $\epsilon_{train}=1.0$ and testing with lower $\epsilon_{test}$ levels of $0.00,0.10,$ and $0.25$, results in a drastic degradation in accuracy for AT for corruption levels greater than 0%. This performance collapse is not observed when using A-RRM. Given that it may be difficult to anticipate the adversarial regime in production environments, A-RRM seems to confer a greater benefit than AT.

### v2

@@ 3.4.1 Loss-reweighting via Data-Driven Wasserstein Formulation
Given a sample $\{(1,x_{i},y_{i})\}_{i=1}^{N}$, just as in Section 3.2, we can opt not to take as granted the resulting empirical distribution $\mathcal{D}_{N}$ because of the possibility that $|C|$-many have incorrect labels (i.e., $y_{i}\neq\tilde{y}_{i})$. Instead, we will admit alternative distributions obtained by shifting the $\mathcal{D}_{N}$’s probability mass off “contaminated” tuples $(1,x_{i},y_{i})_{i\in C}$ to possibly $(0,x_{i},y_{i})$, $(1,x_{i},\tilde{y}_{i})$, or even some other tuple $(1,x_{j},\tilde{y}_{j})$ with $j\notin C$ for example - equivalently, eliminating, correcting, or replacing the sample, respectively. In order to admit such favorable corrections to $\mathcal{D}_{N},$ we can consider the optimistic (Nguyen et al., 2019; Staib & Jegelka, 2017) data-driven problem
@@
in which for each parameter tuning $\theta$, $v_{N}(\theta)$ measures the expected auxiliary loss with respect to the most favorable distribution within an $\epsilon$-prescribed $W_{1}$ (1-Wasserstein) distance of $\mathcal{D}_{N}$. It turns out that a budgeted deviation of the weights alone (and not the feature-label pairs) can approximate (up to an error diminishing in $N$) $v_{N}(\theta)$. More precisely, we derive the following approximation along similar lines to Staib & Jegelka (2017).
@@
$v_{N}^{MIX}(\theta):=\min_{u_{1},\ldots,u_{N}}~{}$ $\sum_{i=1}^{N}(\frac{1}{N}+u_{i})\cdot J(\theta;x_{i},y_{i})+\gamma_{\theta}\lvert\lvert u\rvert\rvert_{1}$
@@
In particular, $-\gamma_{\theta}\leq\min_{i}J(\theta;x_{i},y_{i})$, and any member $i\in\{i:J(\theta;x_{i},y_{i})>\gamma_{\theta}\}$ is down-weighted to zero, i.e., $u_{i}^{*}=-\frac{1}{N}$ for any $u^{*}$ solving $v_{N}^{MIX}(\theta).$
@@
Hence, in some sense, we can say that RRM is an optimistic methodology but that it is less optimistic than the data-driven Wasserstein approach.
3.5 A-RRM/RRM Algorithm
Towards solving problem (3) in the two decisions $\theta$ and $u$, we proceed iteratively with a block-coordinate descent heuristic outlined in Algorithm 1, whereby we update the two separately in cyclical fashion. In other words, we update $\theta$ while holding $u$ fixed, and we update $u$ whilst holding $\theta$ fixed. The update of $\theta$ is an SGD step on a batch of $s$-many samples. The update of $u$ is a linear program, but can also be completed quicker via Corollary 3.1.1.
@@ 3.5 A-RRM/RRM Algorithm
0:  Loss Function $J(\theta;x,y)$, training perturbation $\epsilon\in[0,1]$, Number of epochs $\sigma,$ Batch size $s\geq 1$, learning rate $\eta>0,$ threshold $\gamma>0$, reweighting step $\mu\in(0,1)$. (optional) contamination estimate $C^{\prime}\in[0,1]$.
@@
     $Iter\leftarrow Iter+1$
     $\theta\leftarrow\texttt{GradientStep}(\sigma,s,\eta,\epsilon,\theta,u)$
     if $C^{\prime}$ provided then
        $\ell_{1-C^{\prime}}\leftarrow\max\{\ell:\frac{|\{j:J(\theta;x_{j}^{b},y_{j}^{b})>\ell\}|}{N}\geq C^{\prime}\}$
        $\gamma\leftarrow\ell_{1-C^{\prime}}-\min_{i}J(\theta;x_{i}^{b},y_{i}^{b})$
        $\mu\leftarrow 1$
     end if
     $u\leftarrow$ Re-weight($\gamma,\mu$)
  until Desired Validation Accuracy or Loss, or $Iter$ is sufficiently large
  Return $\theta$
Algorithm 2 GradientStep($\sigma,s,\eta,\epsilon,\theta,u$)
0:  $\sigma\in\mathbb{Z}_{\geq 1},s\in\mathbb{Z}_{\geq 1},u\in U,\eta>0,\epsilon\in[0,1],\theta$
@@
        $\{(x^{b}_{i},y^{b}_{i})\}_{i=1}^{s}\leftarrow\text{draw batch from }\{(x_{i},y_{i})\}_{i=1}^{N}$
@@
  Return $\theta$
Algorithm 3 Re-weight($\gamma,\mu$)
0:  threshold $\gamma>0$, re-weighting step $\mu\in(0,1)$.
@@
  Return $u\leftarrow\mu u^{*}+(1-\mu)$
Apart from the learning rate $\eta$ that is (industry) standard in gradient-based algorithms, Algorithm 1 can be parameter-less; more precisely, the step-size $\mu$ and threshold $\gamma$ parameters can in fact be auto-tuned, in a manner that follows from Corollary 3.1.1, so as to precisely control how many examples are pruned in each re-weighting step, as might be desired or guided by a contamination estimate $C^{\prime}.$
3.5.1 Auto-Tuning: Precise Sample Pruning with a Contamination Estimate $C^{\prime}$
As outlined in Algorithm 1, if an estimate $C^{\prime}$ of the true contamination $C$ is on hand, then upon setting $\gamma=\ell_{1-C^{\prime}}-\min_{i}J(\theta;x_{i},y_{i})$, as detailed in Algorithm 1, where $\ell_{1-C^{\prime}}$ is approximately the $(1-C^{\prime})-$th quantile of the costs, or equivalently, $\max\{J(\theta;x_{i},y_{i}):\frac{\{j:J(\theta;x_{j},y_{j})>J(\theta;x_{i},y_{i})\}}{N}\geq C^{\prime}\}$, then the fraction of observations “pruned” $\frac{|\chi|}{N}=\frac{|\{i:J(\theta;x_{i},y_{i})>\gamma+\min_{j}J(\theta;x_{j},y_{j})\}|}{N}$ is at least $C^{\prime}$, once the hyperparameter $\mu$ is also fixed to $1$. This is essentially the re-weighting action stated in Corollary 3.1.1. We refer the reader to Section 6.1 for experiments on this form of RRM.
3.5.2 A-RRM $(\epsilon>0)$ versus RRM $(\epsilon=0)$
We will refer to Algorithm 1 as A-RRM versus RRM depending on $\epsilon$. The key difference lying in the execution of (Algorithm 2’s) GradientStep.
A-RRM: (Adversarial) Feature Perturbation ($\epsilon>0$) & Label Contamination $C$
When $\epsilon>0$, then GradientStep executes a Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2015) adversarial attack, whereby each time a batch $(x^{b}_{i},y^{b}_{i})_{i=1}^{s}$ is drawn, every $x^{b}_{i}$ is perturbed via $x^{b}_{i}+\epsilon\cdot sign(\nabla_{x}J(\theta,x^{b}_{i},y^{b}_{i}))$. We note that other adversarial perturbation methods can be substituted in place of this choice of attack. Then a Stochastic gradient descent (SGD) step is taken.
RRM: Label Contamination $C$ Only ($\epsilon=0$)
When $\epsilon=0,$ then GradientStep simply reduces to an SGD step.
3.5.3 On Complexity
Each iteration of Algorithm 1 is comprised of two different tasks: 1.) GradientStep and 2.) Re-weight. Task 1’s gradient step calculations are entirely standard practice. Task 2 amounts to the polynomial-sized (in training data) linear program of (4), for which theoretical efficiency is well-established, and commercial solvers abound (we used CPLEX 12.8.0); hence, complexity is poly-time in the training data size, and the scaling of computation with larger datasets should not be limiting in nature. In fact, Theorem 3.1 and Corollary 3.1.1 indicate there exists an exploitable structure for $N\mathrm{log}N$ execution of Task 2. We refer the reader to Sections 6.1 and 6.2 for experiments on MNIST-3 and CIFAR-10 with average computation times of Algorithm 3 reported.
@@ 4 Datasets
The datasets utilized to evaluate RRM possess high-quality labels. So, to accurately assess our method under contaminated training regimes, we perturb the training datasets to achieve various types and levels of contamination $C$. In all cases the test set used for evaluation is nearly (Northcutt et al., 2021) pristine. A detailed description of the datasets utilized is provided below.
MNIST-10(3) (LeCun & Cortes, 2010): A multi-class classification dataset consisting of 70000 images of digits zero through nine. 60000 digits are set aside for training and 10000 for testing. Training labels are swapped for different, randomly selected digits to achieve desired levels of $C$. In addition to the 10-class dataset, we utilize a smaller 3-class version consisting only of digits 0, 1, and 2.
CIFAR-10 (Alex, 2009) A multi-class classification dataset consisting of 60000 32x32 color images of 10 classes, with 6000 images per class. 50000 images are set aside for training and 10000 for testing. As with MNIST, to achieve the desired $C$ training labels are swapped for different, randomly selected image classes.
@@ 5 Architectures
Table 1: CIFAR-10 Test accuracy (%) comparison between CCE, MAE, and MSE loss functions against baseline GCE method across contamination levels. RRM-wrapped accuracy results are in parentheses. Values marked with ($\dagger$) are from (Zhang & Sabuncu, 2018).
Method Contamination $C$
10% 20% 30%
GCE 91.53 $\pm$ 0.25 89.70 $\pm$ $0.11^{\dagger}$ 89.64 $\pm$ 0.17
CCE (+RRM) 89.94 $\pm$ 0.29 (92.20 $\pm$ 0.26) 86.98 $\pm$ $0.44^{\dagger}$ (90.44 $\pm$ 0.31) 81.90 $\pm$ 0.86 (88.49 $\pm$ 0.33)
MAE (+RRM) 92.04 $\pm$ 0.25 (90.10 $\pm$ 3.56) 83.72 $\pm$ $3.84^{\dagger}$ (87.82 $\pm$ 4.07) 88.28 $\pm$ 3.52 (82.98 $\pm$ 6.55)
MSE (+RRM) 91.83 $\pm$ 0.25 (93.04 $\pm$ 0.17) 89.43 $\pm$ 0.37 (91.43 $\pm$ 0.32) 86.34 $\pm$ 0.22 (89.95 $\pm$ 0.55)
We do not strive to develop a novel NN architectures capable of defeating current state-of-the-art (SOA) performance in each data domain. Nor do we focus on developing robust architectures as described in (Song et al., 2020). Rather, we select a reasonable NN architecture and measure model performance with and without the application of RRM. This approach enables us to demonstrate the general superiority of RRM under varied data domains and NN architectures. We discuss the underlying NN architectures that we employ in this section.
MNIST-10: The MNIST dataset has been studied extensively and harnessed to investigate novel machine-learning methods, including CNNs (Baldominos et al., 2019). For experiments utilizing MNIST-10, we adopt a basic CNN architecture with a few convolutional layers. The first layer has a depth of 32, and the next two layers have a depth of 64. Each convolutional layer employs a kernel of size three and the ReLU activation function followed by a max-pooling layer employing a kernal of size 2. The last convolutional layer is connected to a classification head consisting of a 100-unit dense layer with ReLU activation, followed by a 10-unit dense layer with softmax activation. In total, there are 159254 trainable parameters. Categorical cross-entropy is employed for the loss function.
MNIST-3: For experiments utilizing MNIST-3 a simple fully-connected architecture is instead utilized that consists of a few layers. The first two layer consists of 320 units, and the third layer consists of 200 units. Each of these dense layers employs the ReLU activation function. The output layer consists of a 3-unit dense layer with softmax activation. In total, there are 417880 trainable parameters.
CIFAR-10: For comparison purposes, the CIFAR-10 experiments adopt the ResNet-34 architecture utilized in Zhang & Sabuncu (2018), along with their data preprocessing and augmentation scheme. We perform 32x32 random crops following padding with 4 pixels on each side, horizontal random flips, and per-pixel mean subtraction.
@@ 6 Experiments and Results
We perform experiments to see how Algorithm 1 performs under two kinds of regimes: (a) Label Contamination only; (b) Feature Perturbation and Label Contamination.
For regime (a), we evaluated RRM on MNIST-3 and CIFAR-10; we refer the reader to the Appendix (Section C) for further experimentation on Toxic Comments, IMDb and Tissue Necrosis. For regime (b), we evaluated A-RRM on MNIST-10.
6.1 RRM with a Provided Contamination Estimate $C^{\prime}$ of $C$
In this section, we suppose access to a lower bound $C^{\prime}$ on the contamination level $C,$ and leverage this estimate with the auto-tuning of $\mu$ and $\gamma$ as discussed in Section 3.5.1.
6.1.1 Case of Perfect Estimate $C^{\prime}=C$
Implementation Details: Using Tensorflow 2.10 (Abadi et al., 2015), 180 epochs of training are performed using the CIFAR-10 dataset and architecture of Sections 4 and 5 for CCE, MAE and MSE loss functions, and training label contamination $C$ of 10%, 20%, and 30%. For comparison, RRM is executed for the same loss functions and with a provided $C^{\prime}$ equal to $C$ for 12 iterations with $\sigma=15$ epochs per iteration, also for a total of 180 epochs. Stochastic gradient descent (SGD) with a learning rate ($\eta$) of $10.0$, momentum of $0.9$ with Nesterov momentum enabled, and a weight decay of $10^{-5}$ are employed for both sets of experiments.
The results from these experiments are also compared to those achieved by GCE. For $C=20$% results are available directly from Zhang & Sabuncu (2018). For $C=10$% and $C=30$% results are re-computed using a Pytorch (Ansel et al., 2024) implementation (Chou, 2019) of GCE under the authors’ reported hyperparameter settings.
Each experiment is executed 5 times and the average test set accuracy and standard deviation is reported in Table 1, with the RRM results placed in parentheses.
Analysis: We benchmark the benefit of RRM-wrapping against GCE, an NN-training method robust to noisy labeling. As Table 1 indicates, both CCE and MSE methods benefit from an RRM approach, across contamination levels. The benefit of RRM for MAE is less definitive. However, the RRM approach with MSE achieves the best results, even when compared to GCE’s performance. The additional computation time incurred from the Re-weight($\gamma,\mu$) step is an average of 3.88 seconds per execution of every iteration, for a total of 46.56 seconds.
6.1.2 Case of Conservative Estimate $C^{\prime}>C$
We repeated the experimental setup of Section 6.1.1 to evaluate the benefit of RRM for MSE, but now with an estimate $C^{\prime}>C$ provided to RRM.
Analysis: Upon comparing Table 2 with Table 1, we confirm that as the estimate $C^{\prime}$ is made larger and larger than $C,$ conferred benefit by RRM is reduced and is generally not as high as when $C^{\prime}=C$. However, despite this, RRM-wrapping still generally outperforms MSE alone.
Table 2: CIFAR-10 Test accuracy (%) of RRM-wrapped MSE with Conservative Estimate $C^{\prime}$ of $C$ on classification task .
Contamination $C$ 10% 20% 30% $C^{\prime}$ $C+10\%$ 91.73 $\pm$ 0.28 90.60 $\pm$ 0.40 88.47 $\pm$ 0.73 $C+5\%$ 92.55 $\pm$ 0.15 91.29 $\pm$ 0.41 89.43 $\pm$ 0.64
6.2 Enhancing Loss Function Methodologies with RRM
RRM can enhance the performance of common loss functions that do not account for label contamination, including categorical cross-entropy (CCE), mean absolute error (MAE), and mean squared error (MSE). In fact, the following experiments indicate RRM can even enhance those methods devised with robustness to noisy labels, such as early-learning regularization (ELR).
Table 3: MNIST-3 Test accuracy (%) comparison of CCE, MAE, and MSE against ELR across contamination levels. RRM-wrapped accuracy results are in parentheses.
Method Contamination Level
55% 60% 65%
ELR (+ RRM) 98 (99) 97 (98) 82 (87)
CCE (+ RRM) 90 (98) 77 (96) 46 (67)
MAE (+ RRM) 98 (98) 96 (98) 62 (68)
MSE (+ RRM) 90 (98) 74 (97) 45 (87)